Run Plan From Run Coordinator

From Tritium Experiments Group
Jump to: navigation, search

Apr 2 Swing

  • kin9 production

Apr 2 Day

  • Hall Access during Beam Studies

Apr 2 Owl

  • Kin9 Production

Apr 1 Swing

  • Finish up kin7 production
  • LHRS move to kin9
  • Kin 9 Production

Apr 1 Day

  • Kin7 Production

Apr 1 Owl

  • Wrap up kin4 production
  • Move LHRS for kin7
  • kin7 production

Mar 31 Swing

  • kin4 production?

Mar 31 Day

  • kin4 production

Mar 31 Owl

  • kin4 production

Mar 30 Day/Swing

  • Continue boiling (whiteboard)
  • If finished with boiling study:
    • Call RC, we'll move the spectrometer to 21.943deg
    • Take kin4 data (plan tabulated on whiteboard)
      • remember to do carbon hole check and singlefoil PS8=10 raster calibration run first

Mar 30 Owl

  • Continue boiling (whiteboard)

Mar 29 Swing

  • Continue (finish?) boiling check (details on whiteboard)
  • After boiling check, proceed to kin4 on the LHRS (RHRS remains at kin16)
  • Kin4:
    • (The kin4 plan will be diagrammed on the whiteboard as well)
    • theta = 21.934 deg
    • current = 22.5 uA
    • Data:
      • 1H 2.2hrs of beam
      • 2H 1.2hrs of beam
      • 3H 2.2hrs of beam
      • 3He 2.9hrs of beam

Mar 29 Day

  • Continue the boiling check (detailed plan on whiteboard)
    • This is likely to continue through the end of day shift

Feb 05 Swing

  • We start with kin 15 for the left HRS, and keep taking data at kin 16 for the right HRS. Please follow the run plan from the white board when we get beam back.
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • Double check our Shiftworker Jobs and Responsibilities to make sure all activities needed from you are on track.
  • Any question? Call the RC if needed.

Feb 05 Day

  • No beam expected. Only TO needed.

Feb 05 Owl

  • Almost done with kin 13 (left HRS), so we should move this spectrometer to kin 15. We will keep taking data at kin 16 for the right HRS.
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • Double check our Shiftworker Jobs and Responsibilities to make sure all activities needed from you are on track.
  • Any question? Call the RC if needed.

Feb 04 Day/Swing

  • Continue with kin 13 (left HRS) and kin 16 (right HRS)
    • Rounds of run (3He, 3H, 2H) as described on white board (1 hour in each square). Will also take 30min run with carbon and one hour run with empty at the end of the kinematics (follow the white board)
    • We may complete kin 13 during this shift, so we will move the L-HRS to kin 15. Call the RC if you are at about 2 hours from completing the run plan.
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • Double check our Shiftworker Jobs and Responsibilities to make sure all activities needed from you are on track.
  • Any question? Call the RC if needed.

Feb 04 Owl

  • Continue with kin 13 (left HRS) and kin 16 (right HRS)
    • Rounds of run (3He, 3H, 2H) as described on white board (1 hour in each square)
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • Double check our Shiftworker Jobs and Responsibilities to make sure all activities needed from you are on track.
  • Any question? Call the RC if needed.

Feb 03 Owl/Day/Swing

  • Continue with kin 13 (left HRS) and kin 16 (right HRS)
    • Rounds of run (3He, 3H, 2H) as described on white board (1 hour in each square)
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • Double check our Shiftworker Jobs and Responsibilities to make sure all activities needed from you are on track.
  • Any question? Call the RC if needed.

Feb 02 Day/Swing

  • Continue with kin 13 (left HRS) and kin 16 (right HRS)
    • Rounds of run (3He, 3H, 2H) as described on white board (1 hour in each square)
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • Double check our Shiftworker Jobs and Responsibilities to make sure all activities needed from you are on track.
  • Beam is expected to be more stable now with Hall C getting a reasonable beam current. MCC plans to keep running in the current configuration today and for the weekend.
  • Any question? Call the RC if needed.

Feb 02 Owl

  • Continue with kin 13 (left HRS) and kin 16 (right HRS)
    • Rounds of run (3He, 3H, 2H) as described on white board (1 hour in each square)
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • Double check our Shiftworker Jobs and Responsibilities to make sure all activities needed from you are on track.
  • Any question? Call the RC if needed.

Feb 01 Swing

  • Continue with kin 13 (left HRS) and kin 16 (right HRS)
    • Rounds of run (3He, 3H, 2H) as described on white board (1 hour in each square)
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • We got hours of beam unavailable until 6PM while MMC was working to tune higher current to other halls. May still experience high trip rates while theu finish this tuning, but we just keep taking data!
  • Double check our Shiftworker Jobs and Responsibilities to make sure all activities needed from you are on track.
  • Any question? Call the RC if needed.

Feb 01 Day

  • Continue with kin 13 (left HRS) and kin 16 (right HRS)
    • Rounds of run (3He, 3H, 2H) as described on white board (1 hour in each square)
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • The machine condition for Hall A is good, they are trying not to touch our parameters while still working to tune higher current to other halls. May still experience high trip rates, just keep taking data!
  • Double check our Shiftworker Jobs and Responsibilities to make sure all activities needed from you are on track.
  • Any question? Call the RC if needed.

Feb 01 Owl

  • Continue with kin 13 (left HRS) and kin 16 (right HRS)
    • Rounds of run (3He, 3H, 2H) as described on white board (1 hour in each square)
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • MCC trying to have a reliable high current condition to other halls. May experience high trip rates, just keep taking data.
  • Any question? Call the RC if needed.

Jan 31 Swing

  • Continue with kin 13 (left HRS) and kin 16 (right HRS)
    • Rounds of run (3He, 3H, 2H) as described on white board (1 hour in each square)
    • If beam gets unavailable for more than one hours, run carbon hole when beam is back to make sure it is centered. Then go back to continue with "rounds" of (3He, 3H, 2H).
  • MCC trying to deliver higher current to other halls, so we may experience high trip rates.
  • Any question? Call the RC if needed.

Jan 15 Swing / Jan 16 Owl

  • Carry out Kin3 plan (on whiteboard)
  • Kin3 Plan (20.578deg, 3.1 GeV):
    • Carbon Hole center check
    • Dummy, 10uA, 15min
    • Optics, 10uA, 10min
    • Home, 20uA, 10min
    • Carbon, 20uA, 10min
    • Carbon, 20uA, PS8=10 only!, 5min
    • Empty, 20uA, 30min
    • 3He, 20uA, 2hr
    • 1H, 20uA, 50min
    • 2H, 20uA, 50min
    • 3H, 20uA, 1hr 40min
    • Empty, 20uA, 30min
    • 3He, 20uA, 2hr
    • 1H, 20uA, 50min
    • 2H, 20uA, 50min
    • 3H, 20uA, 1hr 40min
  • Call RC for spectrometer move to Kin4!

Jan 15 Owl

  • Continue Kin2 plan from Swing shift (on whiteboard)
  • Call RC if uncertain about anything

Jan 14 Swing

  • Finish Kin1 running (see whiteboard)
  • Angle change
    • Call Jack a little in advance
    • Set Left spectrometer to 19.135deg (TECHS ONLY, Jack will come in)
  • Begin Kin2 (always use PS1=PS2=PS3=1 (unless otherwise specified))
    • Carbon Hole centering check
    • multifoil, 10 uA, 10 minutes
    • carbon, 20 uA, 10 minutes
    • carbon, 20 uA, only PS8=100, 5 minutes
    • Empty Cell, 20uA, 30 minutes
    • 3He Cell, 20uA, 1hr 40min
    • 1H Cell, 20uA, 40min
    • 2H Cell, 20uA, 40min
    • 3H Cell, 1hr 20min
    • Empty Cell, 20uA, 30 minutes
    • 3He Cell, 20uA, 1hr 40min
    • 1H Cell, 20uA, 40min
    • 2H Cell, 20uA, 40min
    • 3H Cell, 1hr 20min

Jan 13 Day

  • Carbon Hole centering (with Meekins adjusting encoder values)
  • Check distance from centered on hole to collimator edge, change only BPMBx (should be nearby in positive BPMx)
  • Raster size fine tuning
  • Harp scan with raster on (clock trigger only)
  • BPM pedestal run (Call in Jason)
  • Bullseye scan (with Jason)
  • Trigger checkout (Florian looking a remote scopes)
  • Prescale/deadtime checks with dummy target
  • DAQ checkout (shifters/experts check online plots)
  • optics data:
    • Single-Foil Carbon, raster off
    • Single-Foil Carbon, raster on
    • Multi-Foil Carbon, raster off
    • Multi-Foil Carbon, raster on
  • If still waiting for gas cell permission:
    • Dummy Target, 10uA, raster on

Jan 12 Swing

  • Hall Access --Completed
  • Collimator edge-finding --completed
  • clock run for BPM sanity check
  • Bogdan's raster target edge-finding sanity check
  • carbon hole finding/centering (see how close we can get, may need to ask for BPMBx != BPMAx in order to avoid collimator)
  • Continue with the Commissioning Plan: [1]
    • Do not skip the bullseye scan (even though we took one harp scan earlier today)
    • Doug will do the Arc energy measurement sometime this weekend. (Don't call him for this in the middle of the night...)
  • If startup procedure is completed, take 5uA runs, 30 min each:
    • Single-Foil Carbon, raster off
    • Single-Foil Carbon, raster on
    • Multi-Foil Carbon, raster off
    • Multi-Foil Carbon, raster on
  • If still waiting for gas cell permission:
    • Dummy Target, 10uA, raster on

Jan 12 Owl


Jan 11 Swing

  • Monitor the target
  • Check that RHRS and LHRS are at 15.577deg and 3.1 GeV
  • Take Cosmics on both arms until beam setup begins
  • coordinate with MCC for 5th pass beam setup!

Dec 16 Owl Shift

  • Finish the Boiling Study
  • ~1am: access to move LHRS to 22deg
  • Run on 3H, 3He, 1H, 2H, and empty cell until Day Shift

Dec 15 Swing Shift

  • See WhiteBoard
  • Boiling Study: [2]


Dec 15 Day Shift

  • Remove Sieve from LHRS
  • Run Dummy with nominal raster at I = 22.5, 15, 10, 5, 2.5 uA -- 2 million events each (~15 minutes @ 2kHz)
  • FADC threshold check
  • Run Multifoil with nominal raster at I = 22.5, 15, 10, 5, 2.5 uA -- 1 million events each (~15 minutes @ 2kHz)
  • Access begins at 1pm https://logbooks.jlab.org/entry/3505097
  • When gas cell running is approved:
    • call RC
    • begin target boiling study [3]

Other Links

Brief but reflects the exact runplan.


Documentation and details. Could be out-of-date. Please always check with run coordinator and student experts.